OpenStudy (anonymous):

A rectangular storage container with an open top is to have a volume of 10m^3. The lenght of the base is twice the width. material for the base costs 10$ per square meter. material for the sides costs 6$ per square meter. What is the cost of the materials for the cheapest such container

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

Hey could you give me just a second while I gather my thoughts

OpenStudy (anonymous):

Ok So start with what is known. The total price will be: \[P=(10lw)+2(6lh)+2(6wh)\] Also, we know that the volume is \[10=lwh\] and that the length is twice the width \[l=2w\]

OpenStudy (anonymous):

ok so substitute l=2w into 10=lwh and you get \[10=2w^2h\] and simplify for h, this will be important later \[h=5/w^2\]

OpenStudy (anonymous):

Now if we take the price equation and substitute l=2w, we get \[P=20w^2+24wh+12wh\] now substitute h=5/w2 and you get \[P=20w^2+180/w\]

OpenStudy (anonymous):

Now comes the tricky part; you can either graph the equation and look for a minimum, or you can take the derivative to find the minimum. I'll take the derivative: \[d/dw(20w^2+180/w)=40w-180/w^2\]

OpenStudy (anonymous):

So P' is equal to \[40w-180/w^2\] Now set the equation equal to zero, so that P' is equal to zero (by definition, this shows the minimum and maximum of the equation)

OpenStudy (anonymous):

I got a zero at \[w=1.651\]

OpenStudy (anonymous):

Now solve for h and you get 5/1.651^2 so h=1.834 and l=3.30193

OpenStudy (anonymous):

putting it into the price equation and you get 54.51+72.68+36.34=$163.54

OpenStudy (anonymous):

\[P=$163.54\]